OpenStudy (anonymous):

simplify the expression (x^2-10x+25)^1/2

OpenStudy (anonymous):

i have (x-5)^2=((x-5)^2)1/2=(x-5) then i do [5, infinity) but the answer is wrong

OpenStudy (goformit100):

Use multinomial theorem.

hartnn (hartnn):

simplified form is x-5

hartnn (hartnn):

u need domain ?

OpenStudy (anonymous):

i have x-5 and for my domain answer i put...

OpenStudy (anonymous):



OpenStudy (anonymous):

or should it be...

hartnn (hartnn):

why ?? why can x not be less than 5 ?

OpenStudy (anonymous):